# Quantum Pufferfish Privacy: A Flexible Privacy Framework for Quantum Systems

### v1

@@ VIII-B Quantum Fairness and QPP
In this subsection, we see that quantum fairness can be formulated as a special case of QPP. This also encourages the design of customized fairness models depending on the application of interest, with the assistance of the QPP framework.
Quantum fairness aims to treat all input states equally, meaning that no pair of input states that are close in some sense (e.g., close in trace distance) should have a large difference in their outcomes when processed by an algorithm [24]. For a quantum decision model $\mathcal{A}=\{\mathcal{E},\{\mathrm{M}_{i}\}_{i\in\mathcal{O}}\}$, where a quantum channel $\mathcal{E}$ is followed by a POVM $\{\mathrm{M}_{i}\}_{i\in\mathcal{O}}$ (i.e., a quantum algorithm described by a quantum to classical channel), quantum fairness is defined in [24] as follows.
@@
Remark 17 (Privacy implying fairness).
Choose $D(\rho\|\sigma)=\left\|\rho-\sigma\right\|_{1}/2$ and $d\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)=\frac{1}{2}\sum_{i}\!\left|\mathrm{Tr}\!\left[\mathrm{M}_{i}\mathcal{E}(\rho-\sigma)\right]\right|$, which are commonly used distance metrics [24]. With the QPP framework
@@
$\Theta$ =$=\{P_{X}:P_{X}\in\mathcal{P}(\mathcal{X})\},$
$\mathcal{M}$ =$=\{\mathrm{M}:0\leq\mathrm{M}\leq\mathrm{I}\},$
if $\mathcal{E}$ satisfies $\varepsilon$-QPP, then the quantum decision model $\mathcal{A}=\{\mathcal{E},\{\mathrm{M}_{i}\}_{i\in\mathcal{O}}\}$ is also fair.
Proposition 14 (Fairness guarantee from QPP).
If $\mathcal{E}$ satisfies $\varepsilon$-QPP with $(\mathcal{S},\mathcal{Q},\Theta,{\mathcal{M}})$ defined above, then $\mathcal{A}=\{\mathcal{E},\{\mathrm{M}_{i}\}_{i\in\mathcal{O}}\}$ is $(\alpha,\sqrt{\varepsilon^{\prime}/2})$-fair, where $\varepsilon^{\prime}=\min\{\varepsilon,\varepsilon^{2}/2\}$.
@@
From Proposition 11, the fact that $\mathcal{E}$ is $\varepsilon$-QPP implies
@@
Then consider the measurement channel that performs the following transformation:
@@ IX Variants of Quantum Pufferfish Privacy Framework
Inspired by variants of the classical DP framework proposed using divergences as privacy metrics [69, 70], and variants of quantum DP [10], we propose variants of quantum PP with the use of generalized quantum divergences [28] (see also [27, Definition 7.13] and recall (1)). According to [43], post-processing is an important property that should be satisfied in any privacy framework. A generalized divergence by its very definition satisfies this requirement for any processing by means of a quantum channel. Thus, quantum generalized divergences by design can be employed for defining privacy metrics.
In this section, we focus on QPP in the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$, where ¯$\bar{\mathcal{M}}\coloneqq\{\mathrm{M}:0\leq\mathrm{M}\leq\mathrm{I}\}$. Here we formulate QPP based on generalized divergences, which we denote by $(\boldsymbol{\mathsf{D}},\varepsilon)$-QPP, where $\boldsymbol{\mathsf{D}}$ is a placeholder for the generalized divergence being used.
@@
$(\boldsymbol{\mathsf{D}},\varepsilon)$-QPP satisfies the following properties, which follow from the properties of the underlying generalized divergence $\boldsymbol{\mathsf{D}}$:
  • Post-processing: Given a mechanism $\mathcal{A}$ satisfying $(\boldsymbol{\mathsf{D}},\varepsilon)$-QPP in the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M})}$, for a quantum channel $\mathcal{N}$, the processed mechanism $\mathcal{N}\circ\mathcal{A}$ also satisfies $(\boldsymbol{\mathsf{D}},\varepsilon)$-QPP in the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$.
  • Convexity: Let $\mathcal{A}_{1},\ldots,\mathcal{A}_{k}$ be $(\boldsymbol{\mathsf{D}},\varepsilon)$-QPP in the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$. Take $I$ as a $k$-ary categorical random variable with parameters $(p_{1},\ldots,p_{k})$. Then the mechanism $\mathcal{A}\coloneqq\mathcal{A}_{I}$ also satisfies $(\boldsymbol{\mathsf{D}},\varepsilon)$-QPP in the same framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$, if $\boldsymbol{\mathsf{D}}$ satisfies the direct-sum property, as defined in [27, Eq. (4.3.7)], from which it follows that $\boldsymbol{\mathsf{D}}$ is jointly convex [27, Proposition 4.15].
  • Parallel composability: If $\mathcal{A}_{i}$ satisfies $(\varepsilon_{i},\delta_{i})$-QPP in $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$ for $i\in\{1,\ldots,k\}$, then the composed mechanism, as defined in Theorem 1, satisfies $\left(\boldsymbol{\mathsf{D}},\sum_{i=1}^{k}\varepsilon_{i}\right)$-QPP in the framework $\left(\mathcal{S},\mathcal{Q}^{(k)},\Theta,\bar{\mathcal{M}}^{k}\right)$, if $\boldsymbol{\mathsf{D}}$ satisfies subadditivity (i.e., if $\boldsymbol{\mathsf{D}}(\rho_{1}\otimes\rho_{2}\|\sigma_{1}\otimes\sigma_{2})\leq\boldsymbol{\mathsf{D}}(\rho_{1}\|\sigma_{1})+\boldsymbol{\mathsf{D}}(\rho_{2}\|\sigma_{2})$ for all states $\rho_{1}$, $\rho_{2}$, $\sigma_{1}$, and $\sigma_{2}$).
@@
The methodologies proposed in Section VII can be used to audit the variants of QPP (based on generalized divergences) as well. In this regard, quantum algorithms and procedures for estimating respective generalized divergence (e.g., Rényi relative entropies) would be useful. This provides more motivation for designing efficient and accurate techniques for estimating them, beyond those already established in [71].
Choosing various generalized quantum divergences, we arrive at a wide spectrum of privacy frameworks. Then we can utilize the tools of quantum information theory, particularly relationships between divergences, to characterize the relative strength of those privacy frameworks. This then provides an operational interpretation of those divergences as privacy metrics.
@@ IX-A Variants Based on Rényi Divergences
First, we define the following quantities. The measured Rényi divergence of order $\alpha\in(0,1)\cup(1,\infty)$ is defined as [72, Eq. (17)]
@@
with $p(x)\coloneqq\mathrm{Tr}[\mathrm{M}_{x}\rho]$ and $q(x)\coloneqq\mathrm{Tr}[\mathrm{M}_{x}\sigma]$ for $\mathcal{M}$ corresponding to a POVM $\{\mathrm{M}_{x}\}_{x\in\mathcal{X}}$. The Rényi preparation divergence of order $\alpha\in(0,1)\cup(1,\infty)$ is defined as [73, 74]
@@
where ${\mathsf{D}}_{\alpha}$ is an arbitrary quantum Rényi divergence that satisfies data processing [75, Eq. (3.7)].
@@
Choosing the preparation divergence $\hat{\mathsf{D}}_{\alpha}=\boldsymbol{\mathsf{D}}$ in $(\boldsymbol{\mathsf{D}},\varepsilon)$-QPP gives the strongest privacy metric that satisfies the post-processing property for the family of quantum Rényi divergences of order $\alpha$, while $\check{\mathsf{D}}_{\alpha}=\boldsymbol{\mathsf{D}}$ gives the weakest privacy metric from that same family of divergences. That is, we have that
@@
The sandwiched Rényi relative entropy in (8) satisfies data processing for $\alpha\in[1/2,1)\cup(1,\infty)$ and the quantum relative entropy in (6) also satisfies data processing. Thus, both of these are candidates for a generalized divergence.
@@
The first implication in both (283) and (286) follows from Proposition 10. Then for the second implication, recall from item 3 of Proposition 2 that $\overline{\mathsf{D}}^{\delta}(\rho\|\sigma)\leq\mathsf{D}_{\max}^{\delta}(\rho\|\sigma)$. Then applying [38, Propositions 5 and 6], we arrive at:
$\overline{\mathsf{D}}^{\delta}(\rho\|\sigma)$ $\leq\widetilde{D}_{\alpha}(\rho\|\sigma)+\frac{1}{\alpha-1}\ln\!\left(\frac{1}{\delta^{2}}\right)+\ln\!\left(\frac{1}{1-\delta^{2}}\right),$ (290)
@@
Note that the chain of implications in (283) holds for every Rényi divergence $\mathsf{D}_{\alpha}$ satisfying data processing, beyond just $\widetilde{\mathsf{D}}_{\alpha}$, because data processing is the key property to adapt Proposition 10 and [38, Proposition 6] (see Eqs. (K51) and (K52) therein).
Remark 22 (Comparison to existing results for QDP and classical DP).
For the specific instantiation of QDP, Lemmas V.4 and V.5 of [10] show that
@@
where the approximation holds for small $\delta$. Our first implication in (283) is tight compared to this since $\varepsilon^{\prime}\leq\varepsilon$ and the second implication is also tight (i.e, $\varepsilon^{\star}\leq\bar{\varepsilon}$) if we choose $\delta^{2}\leq 1-2^{\left(-\frac{1}{\alpha-1}\right)}$ in the small $\delta$ regime.
@@ IX-B Variant Incorporating Entanglement
In this subsection, we propose a variant of QPP that incorporates reference systems, which enables us to observe the effect of entanglement on the privacy of the system of interest. This variant is applicable in identifying whether there is an advantage in designing private systems with the association of other quantum systems.
@@
Since there is no restriction on the reference systems, the supremum in Definition 8 is also taken over the dimension of the reference system $R$. Thus, the supremum here is over an unbounded reference system $R$. However, following from isometric invariance of generalized divergences, together with purification and the Schmidt decomposition, it suffices to take the supremum over pure bipartite states with the reference system $R$ isomorphic to the channel input system $A$. Similar to the QPP framework, and other variants of QPP, the above variant incorporating reference systems also satisfies post-processing, convexity, and parallel composability (similar to Remark 18).
Remark 23 (Adaptive composition with reference systems).
Let $\mathcal{A}_{i}$ be a $(\boldsymbol{\mathsf{D}}^{R},\varepsilon_{i})$-QPP algorithm for $i\in\{1,2\}$. Consider the adaptive composition of $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, as shown in Fig. 10: First $\mathcal{A}_{1}$ is applied on the upper system, then the quantum channel $\mathcal{N}$ acts on both systems, and lastly $\mathcal{A}_{2}$ acts on the upper system. Adaptive composition of $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ also satisfies $(\boldsymbol{\mathsf{D}}^{R},\varepsilon_{1})$-QPP. This shows that the adaptive composition illustrated in Fig. 10 does not degrade privacy, showcasing the strength of the privacy framework in Definition 8.
@@
This follows from the data-processing inequality for the underlying generalized divergence $\boldsymbol{\mathsf{D}}$, choosing the channel as the partial trace over the reference system $R$. Also it indicates that in certain scenarios where the system of interest has entanglement with some other reference systems, general QPP guarantees defined in earlier sections may not be sufficient. It is important to note that the choice of $\boldsymbol{\mathsf{D}}$ could heavily affect the design of useful privacy frameworks with entanglement (recall the example in Remark 3 with the choice $\boldsymbol{\mathsf{D}}=\mathsf{D}_{\max}$, along with the equivalence of $\mathsf{D}_{\max}$ to $\varepsilon$-QPP, as given in (225)).
@@ X Concluding Remarks and Future Directions
In conclusion, our proposed notion of QPP provides a flexible privacy framework for quantum systems. It gives rise to an operational interpretation of the DL divergence, which was used to study properties of QPP mechanisms and characterize privacy-utility tradeoffs. We explored quantum privacy mechanisms (e.g., via the depolarization channel) and presented a methodology for auditing quantum privacy.
Future research directions are abundant and include privacy-utility analysis of specific quantum estimation tasks, designing efficient quantum algorithms that achieve QPP, analysis of the type-II error of our quantum privacy auditing pipeline, tight characterizations of parallel composability of QPP mechanisms, and many more. Another interesting research direction is to devise efficient methods for computing the DL divergence using quantum algorithms, which would enable the auditing of quantum privacy. In the longer term, the proposed framework could lay the foundations for privacy-preserving learning in quantum systems, and we aim in future work to explore various directions in that space.
Acknowledgements—We acknowledge helpful discussions with Rochisha Agarwal, Hansadi Jayamaha, Kaiyuan Ji, Margarite LaBorde, Hemant Mishra, Dhrumil Patel, Aby Philip, Soorya Rethinasamy, and Vishal Singh. We are also indebted to Prof. Çişik Balyk for sharing his valuable insights. TN and MMW acknowledge support from the National Science Foundation under Grant No. 1907615. ZG is partially supported by NSF grants CCF-2046018, DMS-2210368, and CCF-2308446, and the IBM Academic Award.
Appendix A Proof of Lemma 2
The proof given below is closely related to the proof of [65, Lemma 6.21], but there are some subtle differences and so we provide it here for completeness.

### v2

@@ VIII-B Quantum Fairness and QPP
We now demonstrate that quantum fairness can be viewed as a special case of QPP, which should encourage the design of customized fairness models via the QPP framework.
Quantum fairness seeks to treat all input states equally, meaning that all pairs of input states that are close in some metric (e.g., close in trace distance) should yield similar outcomes when processed by an algorithm [34]. For a quantum decision model $\mathcal{A}=\{\mathcal{E},\{\mathrm{M}_{i}\}_{i\in\mathcal{O}}\}$, where a quantum channel $\mathcal{E}$ is followed by a POVM $\{\mathrm{M}_{i}\}_{i\in\mathcal{O}}$ (i.e., a quantum algorithm described by a quantum to classical channel), quantum fairness is defined in [34] as follows.
@@
Proposition 11 (Fairness guarantee from QPP).
Let $D(\rho\|\sigma)=\left\|\rho-\sigma\right\|_{1}/2$ and $d\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)=\frac{1}{2}\sum_{i}\!\left|\mathrm{Tr}\!\left[\mathrm{M}_{i}\mathcal{E}(\rho-\sigma)\right]\right|$. Fix
@@
$\Theta$ $={\mathcal{P}_{2}\big{(}\mathcal{D}(\mathcal{H})\big{)}},$
$\mathcal{M}$ =$=\{\mathrm{M}:0\leq\mathrm{M}\leq\mathrm{I}\}.$
If $\mathcal{E}$ satisfies $\varepsilon$-QPP with $(\mathcal{S},\mathcal{Q},\Theta,{\mathcal{M}})$ above, then $\mathcal{A}=\{\mathcal{E},\{\mathrm{M}_{i}\}_{i\in\mathcal{O}}\}$ is $(\alpha,\sqrt{\varepsilon^{\prime}/2})$-fair, where $\varepsilon^{\prime}=\min\{\varepsilon,\varepsilon^{2}/2\}$.
@@
From Proposition 10, $\mathcal{E}$ being $\varepsilon$-QPP implies
@@
Then, consider the measurement channel that performs the following transformation:
@@ IX Variants of Quantum Pufferfish Privacy Framework
We now propose variants of QPP via generalized divergences (as defined in (1)). [^10] Due to the definition of generalized divergences, the privacy notions defined based on them inherently satisfy post-processing. We provide an operational interpretation of generalized divergences as privacy metrics and characterize the relative strength between them.
Here, we focus on QPP in the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$ and formulate QPP based on generalized divergences, which we denote by $(\boldsymbol{\mathsf{D}},\varepsilon)$-QPP, where $\boldsymbol{\mathsf{D}}$ is a placeholder for the generalized divergence being used.
@@
Indeed, Definition 7 encompasses variants of classical DP [81, 82] and QDP [15], which rely on Rényi divergences as the generalized divergence along with the appropriate choice of QPP framework $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$ stated in Section III-C3 and Section III-C1, respectively.
@@
Post-processing holds, by the definition of generalized divergences. Convexity follows if $\boldsymbol{\mathsf{D}}$ satisfies the direct-sum property, as defined in [37, Eq. (4.3.7)], from which it follows that $\boldsymbol{\mathsf{D}}$ is jointly convex [37, Proposition 4.15].
Parallel composability: If $\mathcal{A}_{i}$ satisfies $(\varepsilon_{i},\delta_{i})$-QPP in $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$ for $i\in\{1,\ldots,k\}$, then the composed mechanism, as defined in Theorem 1, satisfies $\left(\boldsymbol{\mathsf{D}},\sum_{i=1}^{k}\varepsilon_{i}\right)$-QPP in the framework $\left(\mathcal{S},\mathcal{Q}^{(k)},\Theta,\bar{\mathcal{M}}^{k}\right)$, if $\boldsymbol{\mathsf{D}}$ satisfies subadditivity (i.e., if $\boldsymbol{\mathsf{D}}(\rho_{1}\otimes\rho_{2}\|\sigma_{1}\otimes\sigma_{2})\leq\boldsymbol{\mathsf{D}}(\rho_{1}\|\sigma_{1})+\boldsymbol{\mathsf{D}}(\rho_{2}\|\sigma_{2})$ for all states $\rho_{1}$, $\rho_{2}$, $\sigma_{1}$, and $\sigma_{2}$). It is worth noting that by employing this privacy notion based on generalized divergences, we achieve improved composability results, even in scenarios involving joint measurements (recall property 3 of Theorem 2).
@@
The methodologies proposed in Section VII can be used to audit the variants of QPP (based on generalized divergences) as well. In this regard, quantum algorithms and procedures for estimating respective generalized divergences (e.g., Rényi relative entropies) would be useful. This motivates the development of novel techniques for estimating them efficiently and accurately, beyond those already established in [83].
@@ IX-A Variants Based on Rényi Divergences
First, let us recall definitions of the following quantities. The measured Rényi divergence of order $\alpha\in(0,1)\cup(1,\infty)$ is defined as [84, Eqs. (3.116)–(3.117)]
@@
with $p(x)\coloneqq\mathrm{Tr}[\mathrm{M}_{x}\rho]$ and $q(x)\coloneqq\mathrm{Tr}[\mathrm{M}_{x}\sigma]$ for $\mathcal{M}$ corresponding to a POVM $\{\mathrm{M}_{x}\}_{x\in\mathcal{X}}$. The Rényi preparation divergence of order $\alpha\in(0,1)\cup(1,\infty)$ is defined as [85]
@@
where ${\mathsf{D}}_{\alpha}$ is an arbitrary quantum Rényi divergence that satisfies data processing [86, Eq. (3.7)].
@@
Choosing the preparation divergence $\hat{\mathsf{D}}_{\alpha}=\boldsymbol{\mathsf{D}}$ in $(\boldsymbol{\mathsf{D}},\varepsilon)$-QPP gives a stronger privacy metric that satisfies the post-processing property for the family of quantum Rényi divergences of order $\alpha$, while $\check{\mathsf{D}}_{\alpha}=\boldsymbol{\mathsf{D}}$ gives a weaker privacy metric from that same family of divergences. That is, we have that
@@
The above relations follow by the direct application of (222) and Definition 7.
@@
The sandwiched Rényi relative entropy $\widetilde{\mathsf{D}}_{\alpha}$ in (8) satisfies data processing for $\alpha\in[1/2,1)\cup(1,\infty)$ [45, 46], and the quantum relative entropy $\mathsf{D}$ in (6) also satisfies data processing [87]. Thus, both of these are candidates for a generalized divergence.
@@
The first implication in both (230) and (233) follows from Proposition 9. Then for the second implication, recall from item 3 of Proposition 2 that $\overline{\mathsf{D}}^{\delta}(\rho\|\sigma)\leq\mathsf{D}_{\max}^{\delta}(\rho\|\sigma)$. Then applying [48, Propositions 5 and 6], we arrive at:
$\overline{\mathsf{D}}^{\delta}(\rho\|\sigma)$ $\leq\widetilde{\mathsf{D}}_{\alpha}(\rho\|\sigma)+\frac{1}{\alpha-1}\ln\!\left(\frac{1}{\delta^{2}}\right)+\ln\!\left(\frac{1}{1-\delta^{2}}\right),$ (237)
@@
Note that the chain of implications in (230) holds for every Rényi divergence $\mathsf{D}_{\alpha}$ satisfying data processing, beyond just $\widetilde{\mathsf{D}}_{\alpha}$, because data processing is the key property to adapt Proposition 9 and [48, Proposition 6] (see Eqs. (K51) and (K52) therein).
Remark 24 (Comparison to existing results for QDP).
In the special case of QDP, the dependence on the $(\varepsilon,\alpha)$ parameters in (230) provides a strict improvement over previous results. Specifically, Lemmas V.4 and V.5 of [15] show that
@@
where the approximation holds for small $\delta$. Our first implication in (230) is tighter compared to this since $\varepsilon^{\prime}\leq\varepsilon$ and the second implication is also tighter (i.e, $\varepsilon^{\star}\leq\bar{\varepsilon}$) if we choose $\delta^{2}\leq 1-2^{\left(-\frac{1}{\alpha-1}\right)}$ in the small $\delta$ regime.
@@ IX-B Variant Incorporating Entanglement
In this subsection, we introduce a novel variant of QPP that incorporates reference systems. This extension potentially allows us to explore the impact of entanglement on the privacy of the system of interest.
@@
Since there is no restriction on the reference systems, the supremum in Definition 8 is also taken over the dimension of the reference system $R$, which is an unbounded set. However, following from isometric invariance of generalized divergences, together with purification and the Schmidt decomposition, it suffices to take the supremum over pure bipartite states with the reference system $R$ isomorphic to the channel input system $A$.
Remark 25 (Properties).
Similar to variants of QPP (recall Remark 20), this variant incorporating reference systems also satisfies post-processing, convexity, and parallel composability.
Adaptive composition: Let $\mathcal{A}_{i}$ be a $(\boldsymbol{\mathsf{D}}^{R},\varepsilon_{i})$-QPP algorithm for $i\in\{1,2\}$. Consider the adaptive composition of $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, as shown in Fig. 8: First $\mathcal{A}_{1}$ is applied on the upper system, then the quantum channel $\mathcal{N}$ acts on both systems, and lastly $\mathcal{A}_{2}$ acts on the upper system. Adaptive composition of $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ also satisfies $(\boldsymbol{\mathsf{D}}^{R},\varepsilon_{1})$-QPP. This shows that the adaptive composition illustrated in Fig. 8 does not degrade privacy, showcasing the strength of the privacy framework in Definition 8.
@@
This follows from the data-processing inequality for the underlying generalized divergence $\boldsymbol{\mathsf{D}}$, choosing the channel as the partial trace over the reference system $R$. It highlights that in certain scenarios where the system of interest is entangled with other reference systems, the general QPP guarantees defined in earlier sections may not be sufficient.
The choice of $\boldsymbol{\mathsf{D}}$ could heavily affect the design of useful privacy frameworks with entanglement (recall the example in Remark 3 with the choice $\boldsymbol{\mathsf{D}}=\mathsf{D}_{\max}$, along with the equivalence of $\mathsf{D}_{\max}$ to $\varepsilon$-QPP, as given in (198)). Therefore, careful consideration of the appropriate generalized divergence is essential for developing effective and meaningful privacy frameworks that account for entanglement effects.
@@ X Concluding Remarks and Future Directions
This work proposed QPP as a flexible privacy framework for quantum systems. We showed that QPP is captured exactly by the DL divergence, endowing the latter with an operational interpretation of the DL divergence. The DL divergence representation was used to study properties of QPP mechanisms and characterize privacy-utility tradeoffs. As a concrete case study, we explored the depolarization QPP mechanisms and characterized the parameter values to achieve privacy. A methodology for auditing quantum privacy was also developed.
Future research directions are abundant and include privacy-utility analysis of specific quantum estimation tasks, designing efficient quantum algorithms that achieve QPP, analysing the type-II error of our quantum privacy auditing pipeline, providing tight characterizations of parallel composability of QPP mechanisms, and devising efficient methods for computing the DL divergence using quantum algorithms to enable quantum privacy auditing. In the longer term, the proposed framework could lay the foundations for privacy-preserving learning in quantum systems.
Acknowledgements
We acknowledge helpful discussions with Rochisha Agarwal, Hansadi Jayamaha, Kaiyuan Ji, Margarite LaBorde, Hemant Mishra, Dhrumil Patel, Aby Philip, Soorya Rethinasamy, and Vishal Singh. We are also indebted to Prof. Çişik Balyk for sharing his valuable insights.
Appendix A Alternative Proof for Joint-quasi Convexity of DL divergence
By adapting the strong convexity of Hockey-Stick divergence (Proposition II.5 of [15]), with the substitutions $p=q$ and $\gamma_{2}=1$ and $\gamma_{1}=\gamma$ therein, we obtain
$\mathrm{Tr}\!\left[\left(\sum_{i=1}^{k}p_{i}\rho_{i}-\gamma\sum_{i=1}^{k}p_{i}\sigma_{i}\right)_{+}\right]\\ \leq\sum_{i=1}^{k}p_{i}\mathrm{Tr}\!\left[\left(\rho_{i}-\gamma\sigma_{i}\right)_{+}\right]\leq\max_{i}\mathrm{Tr}\!\left[\left(\rho_{i}-\gamma\sigma_{i}\right)_{+}\right].$ (242)
Then by assuming $\max_{i}\mathrm{Tr}\!\left[\left(\rho_{i}-\gamma\sigma_{i}\right)_{+}\right]\leq\delta$, $\gamma$ is a candidate for the optimization of
$\overline{\mathsf{D}}^{\delta}\!\left(\sum_{i=1}^{k}p_{i}\rho_{i}\middle\|\sum_{i=1}^{k}p_{i}\sigma_{i}\right)$. This leads to
$\overline{\mathsf{D}}^{\delta}\!\left(\sum_{i=1}^{k}p_{i}\rho_{i}\middle\|\sum_{i=1}^{k}p_{i}\sigma_{i}\right)\leq\ln(\gamma).$ (243)
The above holds for all $\gamma$ such that $\max_{i}\mathrm{Tr}\!\left[\left(\rho_{i}-\gamma\sigma_{i}\right)_{+}\right]\leq\delta$. Then, optimizing over such $\gamma$, we arrive at joint quasi-convexity:
$\overline{\mathsf{D}}^{\delta}\!\left(\sum_{i=1}^{k}p_{i}\rho_{i}\middle\|\sum_{i=1}^{k}p_{i}\sigma_{i}\right)\leq\max_{i}\overline{\mathsf{D}}^{\delta}\!\left(\rho_{i}\|\sigma_{i}\right).$ (244)
Appendix B Proof of Lemma 2
The proof given below is closely related to the proof of [79, Lemma 6.21], but there are some subtle differences and so we provide it here for completeness.